# RAMAC: Multimodal Risk-Aware Offline Reinforcement Learning and the Role of Behavior Regularization

### v1

@@ 5.2 Qualitative Safety Analysis
We visualize a three representative method’s contrast, risk-aware expressive generator RADAC, risk-neutral expressive generator DiffusionQL, and anchor-perturb risk-averse method ORAAC. Fig. 4 plots the monitored distribution of policies against safe regions. RADAC concentrate probability mass inside or near safe boundary while actively reallocates probability onto high‑return modes that lie within the safety regions. DiffusionQL is tightly concentrated around zero because rare, high penalties depress bootstrapped values near the safe boundary. On the other hand, ORAAC regularizes toward a behavior anchor and thus settles at a low density between‑mode when the anchor lies in risky-region. These features can be further observed in Table 8 in App. D.2
@@ 5.3 Empirical Analysis of Theoretical Insights
We now provide measurements of OOD actions to validate the insights in Sec. 4. For each policy, we report $\varepsilon_{\text{act}}$, the fraction of evaluation actions whose 1-NN distance to the dataset exceeds $\kappa\!\times\!\mathrm{median}\,d_{\mathrm{NN}}$. Sec. 5.2 predicts that (i) (expressive, BC-regularized CVaR objective) should reweight probability within the data manifold, yielding low $\varepsilon_{\text{act}}$. (ii) ORAAC, being less expressive and based on anchor–perturbation, should exhibit higher $\varepsilon_{\text{act}}$ than RADAC; Table 2 confirms this prediction: RADAC retains low-OOD across task, consistent with BC-regularized OOD suppression; ORAAC is consistently higher than RADAC, matching the expected geometric leakage from anchor–perturbation, consistent with the Sec. 4. RADAC achieves risk awareness and expressiveness simultaneously with low-OOD.
Table 2: OOD action rate ($\%\pm$ s.e.) on medium–expert. $\kappa{=}3$).
@@ 6 Conclusions
This paper introduces RAMAC, a model-free framework for risk-aware offline RL using expressive generative policies. This is done by coupling a distributional critic with diffusion/flow actors and simultaneously incorporating a CVaR together with a BC regularization component in the objective function. We provide a theoretical explanation and proof on how BC regularization through the objective can reduce OOD action rate. This observation further provides insight on one possible mechanism through which expressive policies such as diffusion and flow can be effective in the context of offline RL. We confirm our observation through an example and finally, evaluate the method against several baselines on the Stochastic-D4RL benchmark. Our proposed approach improves CVaR on most tasks while maintaining competitive mean return.
Appendix A Proof of Proposition 1
@@
Appendix B Related Works
We review works most relevant to our risk-aware generative trajectory view—policies that map noise to actions through a differentiable path and how safety is enforced therein—while avoiding repetition of the core background already covered in the main text. For a broad taxonomy of offline RL, see Prudencio et al. (2023).
@@ Appendix B Related Works
The main paper reviews diffusion and flow-matching policies (e.g., DiffusionQL, Diffuser, IDQL, FlowQL). Here we note complementary developments not detailed there: (i) DDIM-style imitation learning that accelerates inference while keeping diffusion’s expressiveness (Song et al., 2021a); (ii) real-robot deployments of diffusion policies demonstrating hardware viability (Chi et al., 2023); and on the flow side . These works bolster the case for expressive, differentiable policies but remain risk-neutral in objective design.
Autoregressive generative baselines
Trajectory Transformer (Janner et al., 2021) provides strong risk-neutral baselines by modeling returns/actions autoregressively; diffusion has also been used for open-loop planning (Chen et al., 2023). Because decoding is single-shot, these approaches lack a continuous generative path through which tail-risk gradients can be injected, leading to high mean performance without explicit lower-tail control.
@@
Beyond expectation-oriented objectives, risk-sensitive control formalizes tail-aware criteria via coherent/dynamic risk measures for MDPs (Ruszczyński, 2010). Among coherent measures, CVaR admits sampling- and policy-gradient formulations suitable for RL (Tamar et al., 2015b; a), and has been linked to robustness via CVaR–robust trade-offs (Chow et al., 2015). In the offline regime, safety is often operationalized as high-confidence off-policy evaluation/improvement from fixed logs—e.g., HCOPE/HCPI and SPIBB (Thomas et al., 2015; Laroche et al., 2019)—which bound deployment risk yet do not address how expressive generators should receive lower-tail gradients.
@@
Concurrent actor–critic lines that couple diffusion with value learning remain expectation-oriented: (Anonymous, 2025) stabilizes online diffusion actors with distributional critics and double-$Q$ but does not backpropagate CVaR along the denoising path ;(Fang et al., 2025)formulates offline constrained policy iteration as diffusion noise regression under KL/BC regularization ; and (Ma et al., 2025) studies efficient online diffusion control from an energy-based perspective. Distributional SAC variant (Ma et al., 2020) improve risk sensitivity via value-law estimation—typically with Gaussian policies—yet still lack CVaR shaping ; the diffusion-policy instantiation (Liu et al., 2025) targets multi-modality but likewise reports no CVaR along the multi-step generation . Risk-averse offline methods relying on behavior priors—e.g.(Urpí et al., 2021) ,and diffusion-prior (Chen et al., 2024)—use anchor–perturb/mixing mechanisms , while (Ma et al., 2021) imposes conservative distributional critics (value pessimism) . These approaches either (i) optimize expectation-oriented objectives with expressive generators or (ii) control risk via mixing/pessimism, in contrast to our distributional risk shaping without anchor mixing.
Appendix C Implementation Details
@@
RAMAC employs a reparameterized generative actor $a=\psi_{\theta}(s,z)$ so that gradients from the risk term flow through the entire generative trajectory. RADAC instantiates $\psi_{\theta}$ as a denoising diffusion policy with VP schedule and T=5T{=}5 denoising steps; the score network is an MLP (hidden 256–256, SiLU). RAFMAC instantiates $\psi_{\theta}$ as a deterministic flow–matching ODE solved by Euler with flow_steps K=10K{=}10; the velocity field is an MLP (hidden 256–256, SiLU). For both, the actor objective is $\mathcal{L}_{\pi}=\lambda_{\mathrm{BC}}\mathcal{L}_{\mathrm{BC}}+\eta\,\mathcal{L}_{\mathrm{Risk}}$, where $\mathcal{L}_{\mathrm{BC}}$ is the model’s native BC loss (score matching for diffusion, velocity matching for flow), and $\mathcal{L}_{\mathrm{Risk}}=-\mathbb{E}_{s,a\sim\pi_{\theta}}[\mathrm{CVaR}_{\alpha}(Z_{\phi}(s,a))]$ with $\alpha{=}0.1$.
@@
Both variants share a Double IQN critic trained with the quantile Huber loss ($\kappa{=}1$). Two critics $Z_{\phi_{1}},Z_{\phi_{2}}$ are updated against a min target to curb overestimation; quantiles for TD use $\tau,\tau^{\prime}\!\sim\!\mathcal{U}(0,1)$, while the actor’s CVaR term samples $\tau\!\sim\!U(0,\alpha)$.
For a batch $(s,a,r,s^{\prime})\!\sim\!\mathcal{D}$, we generate the bootstrapping action via the actor:
@@
so that gradients (later used for risk shaping) can flow through the full generative trajectory (short reverse diffusion for RADAC; short ODE flow for RAFMAC). Right after specifying RADAC/RAFMAC actor parameters, we clarify the pre-loss stage for the critic before introducing the final loss. Instead of sampling $\tau$, we use a fixed uniform grid
$\mathcal{T}_{N}\;=\;\Big\{\,\tau_{i}\;=\;\tfrac{i-\tfrac{1}{2}}{N}\,\Big\}_{i=1}^{N}.$ (14)
For CVaR at level $\alpha$, let $m\!=\!\lfloor\alpha N\rfloor$; then
@@
This is equivalent in expectation to drawing $\tau\!\sim\!\mathcal{U}(0,\alpha)$ (cf. Eq. 9) but with lower estimator variance.
We form target quantiles on another grid $\mathcal{T}_{N^{\prime}}$ and define the TD residual
@@
With this pre-loss construction, the final critic objective is exactly the quantile-Huber residual minimization in Eq. 7/17; determinism only replaces stochastic $(\tau,\tau^{\prime})$ by $(\tau_{i},\tau^{\prime}_{j})$ from fixed grids.
The critic minimises the quantile-Huber loss (Dabney et al., 2018; Rowland et al., 2019)
@@
with $\kappa\!=\!1$. Averaging over $N\times N^{\prime}$ quantile pairs yields
@@
Optimising Eq. 17 yields a calibrated estimate of the return law, whose lower tail supplies the CVaR gradients used in Step 2 (Sec. 3.2).
@@
Unless noted, we use Adam for all networks (default $3{\times}10^{-4}$), batch size 256, discount $\gamma{=}0.99$, soft target update $\tau_{\text{target}}{=}0.005$, and no LR decay. RAMAC’s (critic LR, IQN size, $\eta$, gradient–norm clipping, optional QQ–target clipping, etc.) are listed in Table 6.
RAFMAC risk weight tuning
we swept $\eta\in\{1,10,50,100,300,1000\}$ and unified to $\eta{=}1000$ for all datasets; critic settings are fixed ($lr_{\text{critic}}{=}3{\times}10^{-4}$, emb_dim=128=128, n_quantiles=32=32) (Table 6).
@@
Appendix D Additional Experimental Results
D.1 More 2D Synthetic Task Results
@@
On the 2D bandit dataset, three BC models show generator-specific patterns. CVAE-BC collapses topology and places probability in the low‑density gap. Diffusion-BC most faithfully reproduces both the ring and the inner cluster with appropriate thickness. Flow-Matching BC renders draws a sharp ring but allocates less mass to the center and shows edges spread slightly outward.
@@
Epoch-by-epoch samples show how the CVaR term reshapes a diffusion policy under BC. Early iterations spread mass over both modes; by $\sim$200 epochs the policy starts vacating the ring; Between 400 and 800 epochs the ring thins and probability shifts inward, while the central cluster grows; By roughly 950 epochs most mass is at the safe center. The final plot at epoch=1000 is in Fig. 3 Throughout, BC keeps samples on-manifold, so lower-tail risk is reduced without collapsing mode.
More toy results Fig. 7
shows the qualitative pattern is consistent across seeds. CVAE‑QL fills the low‑density gap; Diffusion‑QL and Flow-QL stay on the ring (high mean, higher risk), and anchor–perturb variants (ORAAC family) place samples in the inter-mode region. RADAC reassigns almost all probability to the safe center, whereas RAFMAC leaves a thinner ring. We attribute this to geometry: flow matching transports density via a smooth, near‑invertible ODE field, which preserves shape and favors thinning rather than removing the ring under BC; diffusion uses a stochastic reverse process whose stepwise CVaR guidance can reallocate mass across the low‑density gap.
@@
Figure 7: More toy results. The qualitative pattern is unchanged across seeds; we show another seed. RAFMAC’s thinner ring and inward shift arise from CVaR‑shaped flow transport under BC.
[…]
Environment, Dataset Algorithm Mean CVaR
HalfCheetah-m-e CQL $66.66\!\pm\!13.17$ $135.39\!\pm\!27.71$
CODAC $2\!\pm\!0.16 0.11±0$$1\!\pm\!0.25$
ORAAC $796.06\!\pm\!30.28$ $742.94\!\pm\!22.95$
FlowQL $844.14\!\pm\!16.15$ $754.44\!\pm\!27.26$
DiffusionQL $20.71\!\pm\!18.89$ $76.39\!\pm\!14.39$
RAFMAC $889.56\!\pm\!38.31$ $736.95\!\pm\!102.54$
RADAC $\mathbf{916.64}\!\pm\!35.80$ $\mathbf{805.25}\!\pm\!15.34$
Walker2d-m-e CQL $21.52\!\pm\!8.68$ $64.88\!\pm\!18.32$
CODAC $23.96\!\pm\!10.56$ $43.88\!\pm\!13.28$
ORAAC $969.62\!\pm\!442.36$ $358.55\!\pm\!682.29$
FlowQL $1309.48\!\pm\!233.72$ $468.15\!\pm\!416.61$
DiffusionQL $32.38\!\pm\!64.15$ $116.19\!\pm\!46.39$
RAFMAC $\mathbf{1822.24}\!\pm\!128.36$ $\mathbf{1127.21}\!\pm\!620.95$
RADAC $1708.68\!\pm\!163.19$ $573.22\!\pm\!894.62$
Hopper-m-e CQL $25.87\!\pm\!13.46$ $111.37\!\pm\!51.69$
CODAC $26.59\!\pm\!47.56$ $150.92\!\pm\!42.19$
ORAAC $\mathbf{714.15}\!\pm\!243.57$ $\mathbf{374.63}\!\pm\!326.66$
FlowQL $341.16\!\pm\!75.98$ $8.80\!\pm\!84.57$
DiffusionQL $279.97\!\pm\!215.46$ $872.95\!\pm\!589.90$
RAFMAC $281.24\!\pm\!82.07$ $132.33\!\pm\!183.92$
RADAC $130.74\!\pm\!273.53$ $167.29\!\pm\!107.33$
HalfCheetah-m-r CQL $66.21\!\pm\!11.52$ $127.09\!\pm\!37.10$
CODAC $0.11\!\pm\!0.16$ $1.47\!\pm\!0.53$
ORAAC $18.99\!\pm\!34.67$ $34.09\!\pm\!25.47$
FlowQL $434.33\!\pm\!40.45$ $224.73\!\pm\!146.83$
DiffusionQL $279.95\!\pm\!91.48$ $79.93\!\pm\!110.85$
RAFMAC $449.04\!\pm\!73.84$ $144.73\!\pm\!181.54$
RADAC $\mathbf{525.84}\!\pm\!44.61$ $\mathbf{278.65}\!\pm\!151.27$
Walker2d-m-r CQL $16.90\!\pm\!7.56$ $51.49\!\pm\!14.17$
CODAC $33.59\!\pm\!45.29$ $52.63\!\pm\!42.63$
ORAAC $126.94\!\pm\!178.91$ $203.64\!\pm\!338.87$
FlowQL $411.36\!\pm\!70.84$ $5.08\!\pm\!240.85$
DiffusionQL $96.88\!\pm\!198.31$ $48.14\!\pm\!227.71$
RAFMAC $71.69\!\pm\!241.69$ $\mathbf{530.37}\!\pm\!84.57$
RADAC $\mathbf{615.94}\!\pm\!219.44$ $145.21\!\pm\!39.43$
Hopper-m-r CQL $6.25\!\pm\!20.60 11$$18.70\!\pm\!106.89$
CODAC $47.83\!\pm\!32.01$ $160.08\!\pm\!60.90$
ORAAC $18.00\!\pm\!44.92$ $129.25\!\pm\!108.63$
FlowQL $373.16\!\pm\!109.86$ $62.24\!\pm\!203.02$
DiffusionQL $2.79\!\pm\!12.83$ $51.33\!\pm\!36.90$
RAFMAC $303.44\!\pm\!28.95$ $90.73\!\pm\!93.82$
RADAC $\mathbf{385.58}\!\pm\!55.20$ $\mathbf{-8.16}\!\pm\!92.79$
D.2 Extended Stochastic-D4RL Results

### v2

@@ 5.2 Qualitative Safety Analysis
We visualize a contrast among three representative methods: risk-aware expressive generator RADAC, risk-neutral expressive generator DiffusionQL, and the anchor-perturb risk-averse method ORAAC. Fig. 4 plots the monitored distribution of policies against safe regions. RADAC concentrates probability mass inside or near the risk-free boundary while actively reallocating probability onto high‑return modes that lie within these regions. DiffusionQL is tightly concentrated around zero because rare, high penalties depress bootstrapped values near the risk-free boundary. On the other hand, ORAAC regularizes toward a behavior anchor and thus can settle at a low density between modes when the anchor lies in a risky region. These features can be further observed in App. E.2, which plots Pareto frontiers of mean return versus per-episode safety-violation counts, and in App. E.5 and  E.6, where we show empirical return distributions and risk–return frontiers.
@@ 5.3 Empirical Analysis of Theoretical Insights
We now provide measurements of OOD actions to validate the insights in Sec. 4. For each policy, we report $\varepsilon_{\text{act}}$, the fraction of evaluation actions whose 1-NN distance to the dataset exceeds $\kappa\!\times\!\mathrm{median}\,d_{\mathrm{NN}}$. Sec. 5.2 predicts that (i) the BC-regularized CVaR objective should reweight probability within the data manifold, yielding low $\varepsilon_{\text{act}}$. (ii) ORAAC, being based on anchor–perturbation, should exhibit higher $\varepsilon_{\text{act}}$ than RADAC; Table 2 confirms this prediction: RADAC retains low OOD across tasks, consistent with the prediction in Sec. 4; ORAAC is consistently higher than RADAC, matching the expected geometric leakage from anchor–perturbation. Similar RADAC << ORAAC rankings hold under alternative OOD detectors (App. E.3.3). Overall, RADAC achieves risk awareness and expressiveness simultaneously while maintaining low OOD rates.
Table 2: OOD action rate ($\%\pm$ s.e.) on medium–expert ($\kappa{=}3$).
@@ 6 Conclusions
This paper introduces RAMAC, a model-free framework for risk-aware offline RL using expressive generative policies. This is done by coupling a distributional critic with an expressive generative actor and a simple BC+CVaR objective, instantiated in our main experiments as the diffusion-based RADAC with a flow-based variant in the appendix. Our analysis shows that applying BC directly to the deployed actor yields a forward-KL upper bound on per-state OOD mass and its contribution to CVaR degradation, and that prior-anchored perturbation can still leak into low-density regions on thin or non-convex supports. On Stochastic-D4RL benchmarks with rare hazards, RADAC achieves consistently stronger tail returns (CVaR0.1) with competitive or higher means and lower measured OOD rates than recent SOTA baselines in risk-aware / risk-neutral offline RL.
Appendix A Limitations and Future Directions
Practical remedies include adding a distilled one-step RL head to avoid recursive backprop, and using score- or energy–weighted objectives that reduce full-path backprop for diffusion/flow policies [koirala2025flow; park2025flow]; both are natural directions for future improvement. On the critic side, modeling the return distribution itself with diffusion/flow value networks is a natural extension that may improve tail calibration without sacrificing actor expressiveness [agrawalla2025floq; zhang2025d2]. Our theory is deliberately scoped to proposition-level guidance (e.g., KL–OOD relations) and does not provide finite-sample, high-probability guarantees under function approximation or partial coverage. Strengthening these to non-asymptotic guarantees, exploring alternatives to BC ( ff‑divergences or Wasserstein) to trade off mode‑covering vs. mode‑seeking; extending analysis to dynamic/spectral risk and partial observability, and developing risk‑aware offline‑to‑online fine‑tuning that preserves on‑manifold exploration are all promising directions for future work.
Appendix B Proofs
B.1 Proof of Lemma 4.1
Proof.
By assumption, $\pi_{\text{anch}}(\cdot\mid s)$ has a density $p(a\mid s)$ on $B_{\Phi}(b^{\star})$ with $p(a\mid s)\geq c>0$. Thus
$\delta_{s}(\pi_{\text{anch}})=\int_{O_{s}}p(a\mid s)\,da\;\geq\;\int_{B_{\Phi}(b^{\star})\cap O_{s}}p(a\mid s)\,da\;\geq\;c\cdot\lambda\bigl(B_{\Phi}(b^{\star})\cap O_{s}\bigr)>0.$
The final claim follows as long as the density lower bound $c$ on $B_{\Phi}(b^{\star})$ is maintained during training. ∎
B.1.1 Geometric Intuition Behind Lemma 4.1
Lemma 4.1 formalizes the geometric intuition that when the anchor point $b$ lies near a thin or nonconvex region of the support $\mathcal{B}_{s}$, any perturbation policy constrained within a ball centered at $b$ will unavoidably place some mass outside the dataset support $\mathcal{D}_{s}$. The following points provide an intuitive view:
  • Thin support near $b$: When the local margin $m(b)$ is small, any ball $B_{\Phi}(b)$ with $\Phi\gtrsim m(b)$ necessarily overlaps $O_{s}$, so some residual updates produce OOD actions even if $\|\zeta_{\psi}\|\leq\Phi$.
  • Nonconvex support: If $\mathcal{S}_{G}(s)$ is nonconvex (e.g., a ring/annulus), anchors can sit near holes or concavities, again yielding small $m(b)$ and forcing overlap between $B_{\Phi}(b)$ and $O_{s}$.
  • Gradients pushing off the data surface: The residual $\zeta_{\psi}(s,b)$ is trained to increase $Q$ or CVaR and is not constrained to remain tangent to the manifold, so gradients can push along the normal direction toward the boundary of $B_{\Phi}(b)$, further amplifying leakage when $m(b)\leq\Phi$.
Even when a coherent risk objective (e.g., CVaR) is used to train $\zeta_{\psi}$, Lemma 4.1 implies that a nontrivial fraction of actions remains OOD whenever the behavior support is thin or nonconvex.
B.2 Proof of Proposition 1
@@
Appendix C Related Works
We review works most relevant to our risk-aware generative trajectory view—policies that map noise to actions through a differentiable path and how safety is enforced therein—while avoiding repetition of the core background already covered in the main text. For a broad taxonomy of offline RL, see [prudencio2023survey].
@@ Appendix C Related Works
The main paper reviews diffusion and flow-matching policies (e.g., DiffusionQL, IDQL, FlowQL). Here we note complementary developments not detailed there: (i) Diffusion-policy imitation learning with DDIM-style fast sampling that accelerates inference while retaining diffusion expressiveness in visuomotor or robot control [ankile2024juicer; chi2025diffusion]; (ii) on the flow-matching side, recent generative policies for robot manipulation replace diffusion with flow matching to achieve similar expressiveness with faster, more stable inference [gao2025vita; yan2025maniflow; zhang2025flowpolicy]; (iii) transformer-based trajectory and policy models, including trajectory- and sequence-level generative decision-makers such as Decision Diffuser and Decision Transformer, and hierarchical visuomotor transformers for bimanual or gaze-conditioned manipulation [ajay2022conditional; chen2021decision; chuang2025look; janner2022planning; lee2024interact]. These works bolster the case for expressive generative decision-making models but remain risk-neutral in objective design.
@@
Beyond expectation-oriented objectives, risk-sensitive control formalizes tail-aware criteria via coherent/dynamic risk measures for MDPs [ruszczynski2010]. Among coherent measures, CVaR admits sampling- and policy-gradient formulations suitable for RL [tamar2015coherent; tamar2015optimizing], and has been linked to robustness via CVaR–robust trade-offs [chow2015risk]. In the offline regime, safety is often operationalized as high-confidence off-policy evaluation/improvement from fixed logs [laroche2019safe; thomas2015high], which bound deployment risk yet do not address how expressive generators should receive lower-tail gradients.
Mixture-policy CVaR optimization.
[luo2024simple] propose a mixture policy parameterization for CVaR optimization in online RL, where a risk-neutral policy and an adjustable component are combined to form a risk-averse policy and improve the sample efficiency of CVaR policy gradients. While both their work and ours aim to shape the lower tail of returns, their method uses standard parametric actors and does not address offline data, generative policies, or explicit constraints on out-of-distribution actions. By contrast, RAMAC targets offline risk-sensitive control with a single expressive generative actor (diffusion/flow), a distributional critic, BC+CVaR objectives, per-state OOD constraints, and hazard relabeling.
@@
Concurrent actor–critic lines that couple diffusion with value learning remain expectation-oriented:[zhang2025d2] stabilizes online diffusion actors with distributional critics and double-$Q$ but does not backpropagate CVaR along the denoising path;[fang2024diffusion] formulates offline constrained policy iteration as diffusion noise regression under KL/BC regularization; a distributional SAC variant [ma2020dsac] improves risk sensitivity via value-law estimation and distorted expectations (e.g., CVaR), but operates with standard unimodal Gaussian policies in an online setting, rather than shaping a multimodal diffusion/flow actor under behavior-regularized offline constraints; the diffusion-policy instantiation [liu2025distributional] targets multi-modality but likewise reports no CVaR along the multi-step generation. Risk-averse offline methods relying on behavior priors (e.g. [urpi2021risk]), and diffusion-prior (e.g. [chen2024uncertainty]) use anchor–perturb/mixing mechanisms, while [ma2021conservative] imposes conservative distributional critics (value pessimism). These approaches either (i) optimize expectation-oriented objectives with expressive generators or (ii) control risk via mixing or pessimism, in contrast to our distributional risk shaping without anchor mixing.
Appendix D Implementation Details
@@
RAMAC employs a reparameterized generative actor $a=\psi_{\theta}(s,z)$ (as shown in Eq. 8) so that gradients from the risk term flow through the entire generative trajectory. RADAC instantiates $\psi_{\theta}$ as a denoising diffusion policy with VP schedule and T=5T{=}5 denoising steps; the score network is an MLP (hidden 256–256, SiLU) following [wang2022diffusion]. RAFMAC instantiates $\psi_{\theta}$ as a deterministic flow–matching ODE solved by Euler with flow_steps K=10K{=}10; the velocity field is an MLP (hidden 512–512, SiLU) following [park2025flow]. For both, the actor objective is $\mathcal{L}_{\pi}=\mathcal{L}_{\mathrm{BC}}+\eta\,\mathcal{L}_{\mathrm{Risk}}$ (as shown in Eq. 12), where $\mathcal{L}_{\mathrm{BC}}$ is the model’s native BC loss (score matching for diffusion, velocity matching for flow) as shown in Eq. 11, and $\mathcal{L}_{\mathrm{Risk}}=-\mathbb{E}_{s,a\sim\pi_{\theta}}[\mathrm{CVaR}_{\alpha}(Z_{\phi}(s,a))]$ with $\alpha{=}0.1$ as shown in Eq. 10.
@@
As described in Sec. 3.1, both RAMAC (diffusion actor) and RAFMAC (flow actor) use a Double IQN critic $Z_{\phi}(s,a;\tau)$ that returns the $\tau$-quantile of the return distribution. We instantiate two critics $Z_{\phi_{1}},Z_{\phi_{2}}$ and train them with the quantile Huber loss ($\kappa{=}1$), using a standard double-critic scheme where targets are formed with a minimum over a slowly updated target copy $Z_{\bar{\phi}}$ to curb overestimation.
For a batch $(s,a,r,s^{\prime})\!\sim\!\mathcal{D}$, the bootstrap action is always produced by the same generative actor used in the policy loss:
@@
so that the critic learns return quantiles under the current actor and, in turn, supplies tail-sensitive gradients (via CVaR) to the risk-aware actor update (Step 2, Sec. 3.2). Concretely, one can view the temporal-difference (TD) objective as approximating expectations over $\tau,\tau^{\prime}\!\sim\!\mathcal{U}(0,1)$, while the actor-side CVaR objective uses $\tau\!\sim\!\mathcal{U}(0,\alpha)$.
In implementation, we replace stochastic sampling of $(\tau,\tau^{\prime})$ by fixed uniform grids, which yield the same expectation but with lower estimator variance. Instead of drawing $\tau$ at random, we use
$\mathcal{T}_{N}\;=\;\Big\{\,\tau_{i}\;=\;\tfrac{i-\tfrac{1}{2}}{N}\,\Big\}_{i=1}^{N},$ (16)
and approximate $\mathrm{CVaR}_{\alpha}$ by averaging the lowest-$\alpha$ fraction of quantiles. For CVaR at level $\alpha$, let $m\!=\!\lfloor\alpha N\rfloor$; then
@@
This is equivalent in expectation to drawing $\tau\!\sim\!\mathcal{U}(0,\alpha)$ (Eq. 9), but the deterministic grid substantially reduces variance in both the critic loss and the actor’s CVaR objective.
For the TD targets we use another grid $\mathcal{T}_{N^{\prime}}$ and define the residual for each pair of source and target quantiles as
@@
The critic minimises the standard quantile Huber loss [dabney2018implicit; rowland2019statistics]
@@
with $\kappa\!=\!1$. Averaging over all $N\times N^{\prime}$ quantile pairs yields the final critic objective
@@
Thus Eq. 19 is exactly the compact critic objective in Sec. 3.1, with deterministic grids $(\tau_{i},\tau^{\prime}_{j})$ used in place of stochastic $(\tau,\tau^{\prime})$. Optimising this loss yields a calibrated estimate of the return distribution; its lower tail is then aggregated into $\mathrm{CVaR}_{\alpha}$ to provide the tail-aware gradients used in the RAMAC actor loss (Step 2).
@@
Unless noted, we use Adam for all networks (default $3{\times}10^{-4}$), batch size 256, discount $\gamma{=}0.99$, soft target update $\tau_{\text{target}}{=}0.005$, and no LR decay. RAMAC’s specific hyperparameters (critic LR, IQN size, $\eta$, gradient–norm clipping, optional QQ–target clipping, etc.) are listed in Table 6.
RAFMAC risk-weight tuning
For RAFMAC, we swept $\eta\in\{1,10,50,100,300,1000\}$ and unified to $\eta{=}1000$ for all datasets; critic settings are fixed ($lr_{\text{critic}}{=}3{\times}10^{-4}$, emb_dim=128=128, n_quantiles=32=32) (Table 6).
@@
Appendix E Additional Experimental Results
E.1 More 2D Synthetic Task Results
@@
On the 2D bandit dataset, three BC models show generator-specific patterns. CVAE-BC collapses topology and places probability in the low‑density gap. Diffusion-BC most faithfully reproduces both the ring and the inner cluster with appropriate thickness. Flow-Matching BC draws a sharp ring but allocates less mass to the center and shows edges spread slightly outward. These baselines confirm that a suitably trained generative model can represent the full multimodal support of the dataset.
@@
RADAC starts from the same diffusion generator that is first fit to the behavior data, and then applies CVaR-based policy updates on top of the BC objective. In Fig. 6 we fix the RL weight $\eta$ to the value used in Fig. 3 and visualize how the induced policy evolves over training.
Early in training $\sim$200 epochs, the critic has not yet identified the outer ring as hazardous: high-mean returns on the ring dominate the low-quantile signal, so the CVaR-guided updates still allocate substantial mass to the high-return ring while also covering the safe central cluster. As learning proceeds and low-quantile returns on the ring are observed, CVaR increasingly penalizes ring actions; between about 400 and 800 epochs the ring thins and probability mass shifts inward as the central cluster grows. By roughly 950 epochs most mass concentrates near the safe center, matching the final snapshot in Fig. 3.
This trajectory highlights that RADAC uses the expressive diffusion policy to first capture the multimodal behavior distribution and then reallocate mass toward the high-CVaR central mode once its tail risk is correctly estimated, rather than collapsing due to limited capacity. The explicit BC vs.risk trade-off as the RL weight $\eta$ varies is further analyzed in Fig. 9.
On conservatism and hyperparameters in prior baselines.
The same Risky–Bandit geometry also helps explain why conservative tuning of prior methods (DiffusionQL, FlowQL, ORAAC-style anchor-perturbation) does not eliminate their structural failure modes. As defined in App. F.1, behavior actions lie on a thin, high-mean but heavy-tailed outer ring and a lower-mean but light-tailed central Gaussian cluster. Geometrically, each behavior anchor $b$ lies on a narrow manifold with margin $m(b)$ to the low-density gap and to regions that are unobserved in the dataset. Once a perturbation ball $\mathbb{B}_{\varepsilon}(b)$ has $\varepsilon\gtrsim m(b)$, it necessarily crosses into these off-support regions. In this regime, residual or anchor-perturbation policies are not constrained to stay tangent to the data manifold, so even small residuals can systematically produce OOD actions, which in turn leads to unstable Q-estimates and higher exposure to the trap-heavy ring. This is precisely the behavior illustrated in Fig. 3(f-g).
For DiffusionQL and FlowQL, the RL term is always driven by a risk-neutral Q-function. Sweeping the RL scaling coefficient $\eta$ (as shown in Fig. 7) shows the expected dichotomy: as $\eta\to 0$ the policies revert to BC-like behavior, recovering the multimodal density of the behavior data; as $\eta$ increases, mass increasingly concentrates on the high-mean ring, despite its heavy lower tail. FlowQL’s flow-matching prior retains a bit more mass near the center, but the dominant mode still tracks the risky ring. No choice of $\eta$ can simultaneously prevent this tendency and retain a nontrivial RL component, because the objective is fundamentally risk-neutral.
Fig. 8 shows ORAAC-style anchor-perturbation exhibits an analogous tradeoff. When the risk-distortion weight $\lambda$ is set to zero, the method again collapses to BC on the behavior manifold. As $\lambda$ increases, the residual pushes probability mass away from the BC solution: in the Risky–Bandit geometry this either splits modes or routes probability through low-density regions between them, and for larger $\lambda$ the policy can “jump” toward arbitrary corners of the action space, far from any behavior support (as predicted by the analysis in Sec. 4).
In other words, the observed failures of Fig. 3(c-g) are not artifacts of insufficient hyperparameter tuning but a structural consequence of combining risk-neutral or anchor-perturbation objectives with thin, non-convex data supports. RADAC replaces this with an on-manifold, risk-aware update that reduces tail risk without relying on off-support residuals.
For completeness, we also sweep the RL weight $\eta$ in RADAC on the same Risky–Bandit (Fig. 9). When $\eta\to 0$, the objective reduces to pure BC and the diffusion policy faithfully matches the multimodal behavior distribution, reproducing both the outer ring and the central cluster. As $\eta$ increases, the CVaR-guided term gradually reallocates probability mass away from the high-mean but heavy-tailed ring toward the light-tailed central mode while staying on-support: the ring thins and eventually disappears, leaving a compact cluster around the safe center. Unlike DiffusionQL/FlowQL and ORAAC-style anchor–perturbation, this sweep does not create spurious off-support modes or low-density bridges; instead, it yields a smooth trade-off between ring coverage and tail-risk reduction, consistent with the on-manifold, risk-aware behavior-regularized objective in Eq. 12.
@@
Figure 7: Effect of RL weight $\eta$ in Risky-Bandit. Each panel shows a policy trained on the 2D Risky-Bandit with a different RL scaling coefficient $\eta$. As $\eta\to 0$, especially DiffusionQL reverts to BC-like behavior and recover the multimodal behavior density. As $\eta$ increases, probability mass concentrates on the high-mean outer ring despite its heavy lower tail, illustrating the risk-neutral tendency discussed in Sec. 3 and App. E.
Figure 8: Effect of risk weight $\lambda$ in anchor-perturbation baselines. Each panel visualizes policies on the 2D Risky-Bandit as the risk-distortion weight $\lambda$ is swept. For $\lambda=0$, all ORAAC variants reduce to BC on the behavior manifold. As $\lambda$ increases, residual updates push probability mass away from the BC solution: modes can split or leak through low-density gaps, and for larger $\lambda$ the policy may jump toward regions far from any behavior support, consistent with the analysis in Sec. 4.
Figure 9: Effect of RL weight $\eta$ in RADAC. As $\eta$ increases from 0 (pure BC), probability mass smoothly moves from the risky outer ring to the safe central cluster, without creating off-support modes.
E.2 Extended Stochastic-D4RL Results